# Robustness-Congruent Adversarial Training for Secure Machine Learning Model Updates

### v1

@@ 5 Related Work
We discuss here methodologies that are related to our work. Continual Learning. Our work has ties to the research field of continual learning (CL) [29, 30], which aims to continuously re-train machine-learning models to deal with new classes, tasks, and domains, while only slightly adapting the initial architecture. It has however been shown that such continuous updates render CL techniques sensitive to catastrophic forgetting [31, 32], i.e., to forget previously-learned classes, tasks, or domains, while retaining good performance only on more recent data. Several techniques have been proposed to mitigate this issue, including: replay methods [33, 34], which replay representative samples from past data while re-training models on subsequent tasks; regularization terms [35, 36], which promote solutions with similar weights to those of the given model; and parameter isolation [37, 38], which separates weights attributed to the different task to be learned. However, as also explained in [1], we argue that quantifying the regression of models is not strictly connected to CL, as we are neither considering the inclusion of new tasks to learn nor the adaptation to the evolution of the data distribution. Furthermore, we do not restrict our update policy to maintain the same architecture, but we permit its replacement with a better one in terms of both accuracy and robustness.
Backward Compatibility. Our methodology can be included among the so-called backward-compatible [39, 40, 41] learning approaches, which focus on providing updates of machine learning models that can interchangeably replace previous versions without suffering loss in accuracy. This can be achieved by: (i) learning an invariant representation for newer and past data [41]; using different weights when predicting specific samples [42]; and (iii) estimating which samples should be re-evaluated as their labels might be incorrect. However, differently from RCAT, all these techniques only focus on accuracy, ignoring the side-effects on the regression of robustness, as the dramatic drop of performance we have observed when using PCT.
@@ 6 Conclusions and Future Work
Modern machine-learning systems demand frequent model updates to improve their average performance. To this end, more powerful architectures and additional data are often exploited in such systems to update the current models. However, it has been shown that model updates can induce a perceived regression of accuracy in the end users, as the new model may commit mistakes that the previous one did not make. The corresponding samples misclassified by the new model are referred to as negative flips (NFs). In this work, we show that NFs are not the sole regression that machine-learning models can face. Model updates can indeed cause also a significant regression of adversarial robustness. This means that, even if the average robustness of the updated model is higher, some adversarial examples that were ineffective against the previous model can become effective against the new one. We refer to these samples as robustness negative flips (RNFs). To address this issue, we propose a novel algorithm named RCAT, based on adversarial training, and theoretically show that our methodology provides a statistically-consistent estimator, without affecting the usual convergence rate of $O(\nicefrac{{1}}{{\sqrt{n}}})$. We empirically show the existence of RNF while updating robust image classification models, and compare the performance of our RCAT approach with the adversarially-robust version of PCT, namely PCAT. The results highlight that RCAT better handles the regression of robustness, by reducing the number of RNF and keeping almost the same performance as the previous model.
However, our methodology does not completely prevent the regression of robustness, but only mitigates it. Thus, as future research directions, we plan to improve the proposed approach by studying the effect of different loss functions and regularizers. Furthermore, while we only consider image classification in this work, we envision the presence of RNF also in other domains where machine-learning models require frequent updates, such as malware and spam detection. We plan to extend RCAT accordingly, to tackle the complexity of these non-trivial application domains. To conclude, we firmly believe that this first work can set up a novel line of research, which will educate practitioners to evaluate and mitigate the different types of regression that might be faced when dealing with machine-learning model updates.
@@ Acknowledgments
This research has been partially supported by the Horizon Europe project ELSA (grant agreement no. 101070617); and by SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU. This work has been conducted while Daniele Angioni was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with the University of Cagliari.

### v2

@@ 5 Related Work
We now discuss methodologies that are related to our work. Continual Learning. Our work has ties to the research field of continual learning (CL) [33, 34], which aims to continuously retrain machine-learning models to deal with new classes, tasks, and domains, while only slightly adapting the initial architecture. However, it has been shown that such continuous updates render CL techniques sensitive to catastrophic forgetting [35, 36], i.e., to forget previously-learned classes, tasks, or domains, while retaining good performance only on more recent data. Several techniques have been proposed to mitigate this issue, including: replay methods [37, 38], which replay representative samples from past data while retraining models on subsequent tasks; regularization terms [39, 40], which promote solutions with similar weights to those of the given model; and parameter isolation [41, 42], which separates weights attributed to the different task to be learned. However, as also explained in [1], we argue that quantifying the regression of models is not strictly connected to CL, as we are neither considering the inclusion of new tasks to learn nor the adaptation to the evolution of the data distribution. Furthermore, we do not restrict our update policy to maintain the same architecture, but we permit its replacement with a better one in terms of both accuracy and robustness.
Backward Compatibility. Our methodology can be included among the so-called backward-compatible [43, 44, 45] learning approaches, which focus on providing updates of machine learning models that can interchangeably replace previous versions without suffering loss in accuracy. This can be achieved by: (i) learning an invariant representation for newer and past data [45]; using different weights when predicting specific samples [46]; and (iii) estimating which samples should be re-evaluated as their labels might be incorrect. However, unlike RCAT, all these techniques only focus on accuracy, ignoring the side-effects on the regression of robustness, as the dramatic drop in performance we have observed when using PCT.
@@ 6 Conclusions and Future Work
Modern machine-learning systems demand frequent model updates to improve their average performance. To this end, such systems often exploit more powerful architectures and additional data to update the current models. However, it has been shown that model updates can induce a perceived regression of accuracy in the end users, as the new model may commit mistakes that the previous one did not make. The corresponding samples misclassified by the new model are referred to as negative flips (NFs). In this work, we show that NFs are not the sole regression that machine-learning models can face. Model updates can indeed also cause a significant regression of adversarial robustness. This means that, even if the average robustness of the updated model is higher, some adversarial examples that were not found for certain inputs against the previous model can be found against the new one. We refer to these samples as robustness negative flips (RNFs). To address this issue, we propose a novel algorithm named RCAT, based on adversarial training, and theoretically show that our methodology provides a statistically-consistent estimator, without affecting the usual convergence rate of $O(\nicefrac{{1}}{{\sqrt{n}}})$. We empirically show the existence of RNF while updating robust image classification models, and compare the performance of our RCAT approach with PCAT, i.e., the adversarially-robust version of PCT. The results highlight that RCAT better handles the regression of robustness, by reducing the number of RNF and retaining or even improving the same performance as the previous model.
While we have not considered cases in which the data can also change over time, along with the models, we argue that our methodology can be readily applied also under these more challenging conditions. We will better investigate this aspect in future work, considering different application domains in which data quickly evolves over time, demanding frequent model updates, such as in the case of spam and malware detection. Within this context, we also plan to improve the proposed approach by studying the effect of different loss functions and regularizers, along with the investigation of better model selection methods and the implications of the no-free-lunch theorem [47], in particular, related to the trade-off between accuracy and robustness in non-stationary settings. To conclude, we firmly believe that this first work can set up a novel line of research, which will educate practitioners to evaluate and mitigate the different types of regression that might be faced when dealing with machine-learning model updates.
@@ Acknowledgments
This research has been partially supported by the Horizon Europe projects ELSA (grant agreement no. 101070617) and CoEvolution (grant agreement no. 101168560); by SERICS (PE00000014) and FAIR (PE00000013) under the MUR NRRP funded by the EU-NGEU; and by the EU-NGEU National Sustainable Mobility Center (CN00000023), MUR Decree n. 1033—17/06/2022 (Spoke 10). This work was conducted while Daniele Angioni was enrolled in the Italian National Doctorate on AI run by Sapienza University of Rome in collaboration with the University of Cagliari.